OpenStudy (anonymous):

Solve the following system. 3x + 2y = 18 y= -2/3x + 12

OpenStudy (anonymous):

@abdela25

OpenStudy (mertsj):

Replace the y in the first equation with -2/3x + 12

OpenStudy (abdela25):

solve for the variables?

OpenStudy (anonymous):

Er. I think so. Because I'm guessing the second equation is the answer for y in the first equation.

OpenStudy (anonymous):

But it says on the answer key that it's: (-3.6, 14.4)

OpenStudy (mertsj):

Those are the equations of two lines. You job is to find where the lines cross by finding the x and y values that make both equations true.

OpenStudy (abdela25):

in that case the first part would be x=6-2y/3

OpenStudy (mertsj):

So do what I said. Replace the y in the first equation with -2/3 x + 12

OpenStudy (abdela25):

also in the first part y would equal y=9-3x/2

OpenStudy (mertsj):

\[3x+2(\frac{-2}{3}x+12)=18\]

OpenStudy (mertsj):

\[3x+\frac{-4}{3}x+24=18\]

OpenStudy (mertsj):

Finish solving that equation.

OpenStudy (abdela25):

the second part would be y=12-2x/3 and x would equal x=18-3y/2 basically what @Mertsj is saying hes making it easier